### v1

@@ 2.3 Hilbert Spaces for Meshfree Methods
such that if $u\left(x\right)=\int_{-\infty}^{\infty}e^{-\left|\omega\right|+i\omega x}a\left(\omega\right)\,\text{d}\omega$ and $v\left(x\right)=\int_{-\infty}^{\infty}e^{-\left|\omega\right|+i\omega x}b\left(\omega\right)\,\text{d}\omega$, then the inner product on $\mathcal{H}$ is given by $\left(u,v\right)_{\mathcal{H}}:=\left(a,b\right)_{L^{2}\left(\mathbb{R}\right)}.$ Now consider the interpolation problem on $\left\{x_{j}\right\}_{j=1}^{\tilde{N}}\subset S\subset\mathbb{R}$.
@@
That is, $\mathcal{L}^{*}\hat{\boldsymbol{e}}_{j}$ corresponds to the RBF $\phi$ such that $\phi\left(x-x_{j}\right):=4/\left(4+\left(x-x_{j}\right)^{2}\right)$: an inverse quadratic RBF. Indeed, the RBF interpolation matrix $\boldsymbol{\Phi}$ for this problem is given by $\Phi_{jk}=4/\left(4+\left(x_{j}-x_{k}\right)^{2}\right)=\left(\mathcal{L}\mathcal{L}^{*}\right)_{jk}.$
@@
Consider placing a domain $S$ inside a box $\Omega\in\mathbb{R}^{m}$. We consider the set of periodic Fourier basis functions on $\Omega$: $\left\{\phi_{n}\right\}_{n=1}^{\infty}$ where $\phi_{n}\left(\boldsymbol{x}\right)=e^{i\boldsymbol{\omega}_{n}\cdot\boldsymbol{x}}$ for frequencies $\boldsymbol{\omega}_{n}$ such that $\phi_{n}$ is periodic on $\Omega$. Then, define a Hilbert space $\mathcal{H}$ by
@@
where $d=\left\{d_{n}\right\}_{n=1}^{\infty}>0$ is a series that must be chosen. In the case that $\mathcal{F}$ and $\mathcal{G}$ from Subsection 2.1 are linear differential operators of at most order $p$ with bounded coefficient functions, the condition for $\mathcal{L}$ to be bounded (and therefore the method to converge) is that $\left\|\boldsymbol{\omega}\right\|_{2}^{p}d^{-\frac{1}{2}}\in\ell^{2}$, where $\left\|\boldsymbol{\omega}\right\|_{2}:=\left\{\left\|\boldsymbol{\omega}_{n}\right\|_{2}\right\}_{n=1}^{\infty}$ is a sequence of Euclidean norms of the Fourier frequencies (see [31] for further discussion). Another possible construction of this Hilbert space is
@@
In this construction, it is clear that choosing $d_{n}=\left(1+\left\|\boldsymbol{\omega}_{n}\right\|_{2}^{2}\right)^{p}$ corresponds to minimum Sobolev norm interpolation (see [4], for example). The inner product is
@@
This is useful because, in practice, we will always be using a truncated Fourier series. When $d_{n}$ is separable, we can evaluate $\left(\mathcal{L}^{*}\hat{\boldsymbol{e}}_{j}\right)\left(x,y\right)$ with $N_{\omega}=N_{\omega,x}N_{\omega,y}$ Fourier basis functions in $\mathcal{O}\left(N_{\omega,x}+N_{\omega,y}\right)$ operations, rather than $\mathcal{O}\left(N_{\omega,x}N_{\omega,y}\right)$. This means we can typically form the $\boldsymbol{\Phi}=\boldsymbol{VV}^{*}$ matrix in $\mathcal{O}\left(N_{\omega}^{\frac{1}{m}}\left(\tilde{N}+\tilde{N}_{\partial}\right)^{2}\right)$ time for a series in $m$ dimensions rather than the $\mathcal{O}\left(N_{\omega}\left(\tilde{N}+\tilde{N}_{\partial}\right)^{2}\right)$ of computing $\boldsymbol{VV}^{*}$ directly. However, the $\boldsymbol{\Phi}$ matrix is poorly conditioned, so this approach is typically limited to 5 or 6 digits of accuracy.
@@
where $\boldsymbol{V}$ is the matrix such that $\boldsymbol{Va}\in\mathbb{C}^{\tilde{N}+\tilde{N}_{\partial}}$ is a vector consisting of the values of $\mathcal{F}\left(\sum_{n=1}^{N_{\omega}}a_{n}d_{n}^{-\frac{1}{2}}\phi_{n}\right)$ and $\mathcal{G}\left(\sum_{n=1}^{N_{\omega}}a_{n}d_{n}^{-\frac{1}{2}}\phi_{n}\right)$ on $\left\{\boldsymbol{x}_{j}\right\}_{j=1}^{\tilde{N}}$ and $\left\{\boldsymbol{y}_{j}\right\}_{j=1}^{\tilde{N}_{\partial}}$, respectively. In this case, the matrix $\boldsymbol{\Phi}$ for RBF methods corresponds to $\boldsymbol{VV}^{*}$; one could solve $\boldsymbol{VV}^{*}\boldsymbol{\beta}=\boldsymbol{f}$, then the solution to (16) is given by $\boldsymbol{a}=\boldsymbol{V}^{*}\boldsymbol{\beta}$. As stated previously, this approach is poorly conditioned, however, and superior results can generally be obtained from solving (16) using a singular value decomposition or complete orthogonal decomposition. That said, if speed while using a large number of Fourier basis functions is the main concern, forming $\boldsymbol{VV}^{*}=\boldsymbol{\Phi}$ using a separable $d$ as described around Eq. (15) will be less computationally intensive than solving (16) directly. Solving the positive definite system $\boldsymbol{\Phi}\boldsymbol{\beta}=\boldsymbol{f}$ requires $\mathcal{O}\left(\left(\tilde{N}+\tilde{N}_{\partial}\right)^{3}\right)$ operations using a Cholesky decomposition or another direct algorithm, while the cost of forming a decomposition of $\boldsymbol{V}$ is $\mathcal{O}\left(N_{\omega}\left(\tilde{N}+\tilde{N}_{\partial}\right)^{2}\right)$; $N_{\omega}$ may be significantly larger than $\tilde{N}+\tilde{N}_{\partial}$ to ensure $\boldsymbol{V}$ is full rank. See [31] for further discussion on using these Hilbert spaces for solving surface PDEs.
@@ 3 Methods and Analysis
We now present two methods for estimating $\int_{S}f$ on irregular flat domains and surfaces. Throughout this section, we assume that $\mathcal{H}$ is a Hilbert space associated with a particular order of convergence for solving a Poisson problem using a constrained optimization approach (5). Specifically, we assume that there exist constants $p,A_{p},B_{p}>0$ such that if $u\in\mathcal{H}$ and $\Delta_{S}u-f$ has scattered zeros on a set of points $\left\{\boldsymbol{x}_{j}\right\}_{j=1}^{\tilde{N}}\subset S$ with fill distance $h_{\text{max}}$, then $\left\|\Delta_{S}u-f\right\|_{L^{\infty}\left(S\right)}\leq A_{p}h_{\text{max}}^{p}\left\|u\right\|_{\mathcal{H}}$ for a sufficiently smooth $f$ and small enough $h_{\text{max}}$. Furthermore, if $\hat{\boldsymbol{n}}_{\partial S}\cdot\nabla_{S}u=0$ on a set of scattered points $\left\{\boldsymbol{y}_{j}\right\}_{j=1}^{\tilde{N}}\subset\partial S$ with fill distance $h_{\text{max}}$, then $\left\|\hat{\boldsymbol{n}}_{\partial S}\cdot\nabla_{S}u\right\|_{L^{\infty}\left(\partial S\right)}\leq B_{p}h_{\text{max}}^{p}\left\|u\right\|_{\mathcal{H}}$. See the discussion leading up to (12) for an explanation of these assumptions.
@@ 3.1 Method 1: Poisson Solvability
where $\hat{\boldsymbol{n}}_{S}$ is the normal vector to $S$ and $c$ is some constant. If $S$ is a closed surface, the boundary condition is simply removed. The problem (17) is only solvable when $\int_{S}\left(f-cg\right)=0$, or equivalently, when
@@
We can then discretize problem (17). Let $\mathcal{H}$ be a Hilbert space satisfying the assumptions at the beginning of this section. Also assume that $f$ and $g$ are smooth enough so that the solution to (17) for the value of $c$ given by (18) can be extended to some $u_{c_{*}}\in\mathcal{H}$. Let $\left\{\boldsymbol{x}_{j}\right\}_{j=1}^{\infty}$ be dense in $S$, and let $\left\{\boldsymbol{y}_{j}\right\}_{j=1}^{\infty}$ be dense in $\partial S$. Then consider the problem:
@@
Now, this will only be bounded as $\tilde{N},\tilde{N}_{\partial}\to\infty$ when $\int_{S}\left(f-cg\right)=0$ using Proposition 2.1. Then, as $\tilde{N},\tilde{N}_{\partial}\to\infty$, the minimizer $\left(\boldsymbol{f}-c\boldsymbol{g}\right)^{*}\boldsymbol{\Phi}^{-1}\left(\boldsymbol{f}-c\boldsymbol{g}\right)$ as a function of $c$ must approach the value of $c$ where $\int_{S}\left(f-cg\right)=0$. This minimum (assuming $\boldsymbol{\Phi}$ is real) is at
(20) $\tilde{c}_{\tilde{N}}:=\frac{\boldsymbol{g}^{*}\boldsymbol{\Phi}^{-1}\boldsymbol{f}}{\boldsymbol{g}^{*}\boldsymbol{\Phi}^{-1}\boldsymbol{g}},$
where we note that $\boldsymbol{\Phi}^{-1}$ is symmetric. Therefore, as $\tilde{N},\tilde{N}_{\partial}\to\infty$,
@@
So, we have a linear operator to estimate $\int_{S}f$, with real weights as long as $\boldsymbol{\Phi}^{-1}$ is real (which is typical for a reasonable choice of $\mathcal{H}$). To be specific, let $\boldsymbol{w}^{*}=\int_{S}g\boldsymbol{\boldsymbol{g}}^{*}\boldsymbol{\Phi}^{-1}/\left(\boldsymbol{g}^{*}\boldsymbol{\Phi}^{-1}\boldsymbol{g}\right)$, then $\boldsymbol{w}^{*}\boldsymbol{f}\to\int_{S}f$; $\boldsymbol{w}^{*}$ gives integration weights.
We may also want similar expressions to (20) and (21) in terms of inner products in $\mathcal{H}$. Let $v_{g}^{\left(\tilde{N}\right)}$ and $v_{f}^{\left(\tilde{N}\right)}$ be the solutions to
@@
where $F=g$ and $F=f$ to produce $v_{g}^{\left(\tilde{N}\right)}$ and $v_{f}^{\left(\tilde{N}\right)}$, respectively. These may correspond to problems on the domain $S$ that are not actually solvable, but the discretized version will still admit a unique solution. Then, as $\tilde{N},\tilde{N}_{\partial}\to\infty$,
@@ 3.1.1 Analysis
We now study the convergence rate of this method; in particular, we show that it inherits the convergence rate of the underlying meshfree PDE method. Again assuming that fcgf-cg is sufficiently smooth and $c$ is in any bounded interval, let $u_{c}^{\left(\tilde{N}\right)}$ be the solution to (19) for a value of $c$ and $\tilde{N}$, then
@@
It is important to note again here that $\left\|u_{c}^{\left(\tilde{N}\right)}\right\|_{\mathcal{H}}$ may diverge as $\tilde{N}\to\infty$ and $h_{\text{max}}\to 0$; in fact it must diverge if (17) is not solvable for some choice of $c$ by Proposition 2.1. Then, where $\left|\partial S\right|:=\int_{\partial S}1$,
(24) $\left|\int_{S}\Delta_{S}u_{c}^{\left(\tilde{N}\right)}\right|=\left|\int_{\partial S}\hat{\boldsymbol{n}}_{\partial S}\cdot\nabla_{S}u\right|\leq\left|\partial S\right|B_{p}h_{\text{max}}^{p}\left\|u_{c}^{\left(\tilde{N}\right)}\right\|_{\mathcal{H}}$.
@@
In the case that $c=\int_{S}f/\int_{S}g$ and the PDE is solvable, this tells us nothing. However, this holds even when the PDE is not solvable, which gives us an estimate for the rate at which $\left\|u^{\left(\tilde{N}\right)}\right\|_{\mathcal{H}}$ diverges:
@@
Now assume that an exact, extended, strong-form solution $u_{c_{*}}\in\mathcal{H}$ for $c_{*}=\int_{S}f/\int_{S}g$ exists for problem (17). In this case, $\left\|u_{c_{*}}^{\left(\tilde{N}\right)}\right\|_{\mathcal{H}}\leq\left\|u_{c_{*}}\right\|_{\mathcal{H}}$ since $u_{c_{*}}$ is feasible for (19). Recall that $\tilde{c}_{\tilde{N}}=\left(\boldsymbol{g}^{*}\boldsymbol{\Phi}^{-1}\boldsymbol{f}\right)/\left(\boldsymbol{g}^{*}\boldsymbol{\Phi}^{-1}\boldsymbol{g}\right)$ is the minimizer of $\left\|u_{c}^{\left(\tilde{N}\right)}\right\|_{\mathcal{H}}$ as a function of $c$ and that we already know that $\tilde{c}_{\tilde{N}}\to c_{*}$ from Eq. (21). Then,
@@
This estimate depends only on the norm of the PDE solution for $c=c_{*}$ and tells us that convergence of the integral estimate is at the rate at which $\Delta_{S}u_{c_{*}}^{\left(\tilde{N}\right)}$ and $\hat{\boldsymbol{n}}_{\partial S}\cdot\nabla_{S}u$ converge when solving the PDE for $c=c_{*}$.

### v2

@@ 2.3 Hilbert Spaces for Meshfree Methods
such that if $u\left(x\right)=\int_{-\infty}^{\infty}e^{-\left|\omega\right|+i\omega x}a\left(\omega\right)\,\text{d}\omega$ and $v\left(x\right)=\int_{-\infty}^{\infty}e^{-\left|\omega\right|+i\omega x}b\left(\omega\right)\,\text{d}\omega$, then the inner product on $\mathcal{H}$ is given by $\left(u,v\right)_{\mathcal{H}}:=\left(a,b\right)_{L^{2}\left(\mathbb{R}\right)}.$ Now consider the interpolation problem on $\{x_{j}\}_{j=1}^{\tilde{N}}\subset S\subset\mathbb{R}$.
@@
That is, $\mathcal{L}^{*}\hat{\boldsymbol{e}}_{j}$ corresponds to the RBF $\phi$ such that $\phi\,(x-x_{j}):=4/(4+(x-x_{j})^{2})$: an inverse quadratic RBF. Indeed, the RBF interpolation matrix $\boldsymbol{\Phi}$ for this problem is given by $\Phi_{jk}=4/(4+(x_{j}-x_{k})^{2})=(\mathcal{L}\mathcal{L}^{*})_{jk}.$
@@
Consider placing a domain $S$ inside a box $\Omega\subset\mathbb{R}^{m}$. We consider the set of periodic Fourier basis functions on $\Omega$: $\left\{\phi_{n}\right\}_{n=1}^{\infty}$ where $\phi_{n}\left(\boldsymbol{x}\right)=e^{i\boldsymbol{\omega}_{n}\cdot\boldsymbol{x}}$ for frequencies $\boldsymbol{\omega}_{n}$ such that $\phi_{n}$ is periodic on $\Omega$. Then, define a Hilbert space $\mathcal{H}$ by
@@
where $d=\left\{d_{n}\right\}_{n=1}^{\infty}>0$ is a sequence that must be chosen. In the case that $\mathcal{F}$ and $\mathcal{G}$ from Subsection 2.1 are linear differential operators of at most order $p$ with bounded coefficient functions, the condition for $\mathcal{L}$ to be bounded (and therefore the method to converge) is that $\left\|\boldsymbol{\omega}\right\|_{2}^{p}d^{-\frac{1}{2}}:=\{\left\|\boldsymbol{\omega}_{n}\right\|_{2}^{p}d_{n}^{-\frac{1}{2}}\}_{n=1}^{\infty}\in\ell^{2}$, where $\{\boldsymbol{\omega}\}_{2}:=\{\|\boldsymbol{\omega}_{n}\|_{2}\}_{n=1}^{\infty}$ is a sequence of Euclidean norms of the Fourier frequencies (see [36] for further discussion). Another possible construction of this Hilbert space is
@@
In this construction, it is clear that choosing $d_{n}=(1+\|\boldsymbol{\omega}_{n}\|_{2}^{2})^{p}$ corresponds to minimum Sobolev norm interpolation (see [5], for example). The inner product is
@@
This is useful because, in practice, we will always be using a truncated Fourier series. When $d_{n}$ is separable, we can evaluate $(\mathcal{L}^{*}\hat{\boldsymbol{e}}_{j})\left(x,y\right)$ with $N_{\omega}=N_{\omega,x}N_{\omega,y}$ Fourier basis functions in $\mathcal{O}\left(N_{\omega,x}+N_{\omega,y}\right)$ operations, rather than $\mathcal{O}\left(N_{\omega,x}N_{\omega,y}\right)$. This means we can typically form the $\boldsymbol{\Phi}=\boldsymbol{VV}^{*}$ matrix in $\mathcal{O}\,(N_{\omega}^{1/m}(\tilde{N}+\tilde{N}_{\partial})^{2})$ time for a series in $m$ dimensions rather than the $\mathcal{O}\,(N_{\omega}(\tilde{N}+\tilde{N}_{\partial})^{2})$ of computing $\boldsymbol{VV}^{*}$ directly. However, the $\boldsymbol{\Phi}$ matrix is poorly conditioned, so this approach is typically limited to 5 or 6 digits of accuracy.
@@
where $\boldsymbol{V}$ is the matrix such that $\boldsymbol{Va}\in\mathbb{C}^{\tilde{N}+\tilde{N}_{\partial}}$ is a vector consisting of the values of $\mathcal{F}\,(\sum_{n=1}^{N_{\omega}}a_{n}d_{n}^{-1/2}\phi_{n})$ and $\mathcal{G}\,(\sum_{n=1}^{N_{\omega}}a_{n}d_{n}^{-1/2}\phi_{n})$ on $\{\boldsymbol{x}_{j}\}_{j=1}^{\tilde{N}}$ and $\{\boldsymbol{y}_{j}\}_{j=1}^{\tilde{N}_{\partial}}$, respectively. In this case, the matrix $\boldsymbol{\Phi}$ for RBF methods corresponds to $\boldsymbol{VV}^{*}$; one could solve $\boldsymbol{VV}^{*}\boldsymbol{\beta}=\boldsymbol{f}$, then the solution to (16) is given by $\boldsymbol{a}=\boldsymbol{V}^{*}\boldsymbol{\beta}$. As stated previously, this approach is poorly conditioned, however, and superior results can generally be obtained from solving (16) using a singular value decomposition or complete orthogonal decomposition. That said, if speed while using a large number of Fourier basis functions is the main concern, forming $\boldsymbol{VV}^{*}=\boldsymbol{\Phi}$ using a separable $d$ as described around Eq. (15) will be less computationally intensive than solving (16) directly. Solving the positive definite system $\boldsymbol{\Phi}\boldsymbol{\beta}=\boldsymbol{f}$ requires $\mathcal{O}\,((\tilde{N}+\tilde{N}_{\partial})^{3})$ operations using a Cholesky decomposition or another direct algorithm, while the cost of forming a decomposition of $\boldsymbol{V}$ is $\mathcal{O}\,(N_{\omega}(\tilde{N}+\tilde{N}_{\partial})^{2})$; $N_{\omega}$ may be significantly larger than $\tilde{N}+\tilde{N}_{\partial}$ to ensure $\boldsymbol{V}$ is full rank. See [36] for further discussion on using these Hilbert spaces for solving surface PDEs.
@@ 3 Methods and Analysis
We now present two methods for estimating $\int_{S}f$ on irregular flat domains and surfaces. Throughout this section, we assume that $\mathcal{H}$ is a Hilbert space associated with a particular order of convergence for solving a Poisson problem using a constrained optimization approach (5). Specifically, we assume that there exist constants $p,A_{p},B_{p}>0$ such that if $u\in\mathcal{H}$ and $\Delta_{S}u-f$ has scattered zeros on a set of points $\{\boldsymbol{x}_{j}\}_{j=1}^{\tilde{N}}\subset S$ with fill distance $h_{\text{max}}$, then $\|\Delta_{S}u-f\|_{L^{\infty}\left(S\right)}\leq A_{p}h_{\text{max}}^{p}\left\|u\right\|_{\mathcal{H}}$ for a sufficiently smooth $f$ and small enough $h_{\text{max}}$. Furthermore, if $\hat{\boldsymbol{n}}_{\partial S}\cdot\nabla_{S}u=0$ on a set of scattered points $\{\boldsymbol{y}_{j}\}_{j=1}^{\tilde{N}}\subset\partial S$ with fill distance $h_{\text{max}}$, then $\left\|\hat{\boldsymbol{n}}_{\partial S}\cdot\nabla_{S}u\right\|_{L^{\infty}\left(\partial S\right)}\leq B_{p}h_{\text{max}}^{p}\left\|u\right\|_{\mathcal{H}}$. See the discussion leading up to (12) for an explanation of these assumptions.
@@ 3.1 Method 1: Poisson Solvability
where $\hat{\boldsymbol{n}}_{\partial S}$ is the conormal vector along $\partial S$, so that it is tangential to $S$ and normal to $\partial S$, and $c$ is some constant. If $S$ is a closed surface, the boundary condition is simply removed. The problem (17) is only solvable when $\int_{S}\left(f-cg\right)=0$, or equivalently, when
@@
We can then discretize problem (17). Let $\mathcal{H}$ be a Hilbert space satisfying the assumptions at the beginning of this section. Also assume that $f$, $g$, and $S$ are smooth enough, and $\partial S$ is such that a solution to (17) for the value of $c$ given by (18) can be extended to some $u_{c_{*}}\in\mathcal{H}$. For example, in flat domains, for $f,g\in H^{p}(S)$ and a $C^{p+2}$ boundary $\partial S$, an extended solution can be found for $\mathcal{H}=H^{p+2}(\Omega)$ (see Ch. 4, Section 2, Thm. 4 of [20] for PDE regularity, and Thm. A.4 of [19] for the extension). Let $\{\boldsymbol{x}_{j}\}_{j=1}^{\infty}$ be dense in $S$, and let $\{\boldsymbol{y}_{j}\}_{j=1}^{\infty}$ be dense in $\partial S$. Then consider the problem:
@@
Now, this will only be bounded as $\tilde{N},\tilde{N}_{\partial}\to\infty$ when $\int_{S}\left(f-cg\right)=0$ using Proposition 2.1. Noticing that the expression for $\left\|\tilde{u}\right\|_{\mathcal{H}}^{2}$ is simply an upward parabola in $c$, then, as $\tilde{N},\tilde{N}_{\partial}\to\infty$, the minimizer of $\|\tilde{u}\|_{\mathcal{H}}^{2}=(\boldsymbol{f}-c\boldsymbol{g})^{*}\boldsymbol{\Phi}^{-1}(\boldsymbol{f}-c\boldsymbol{g})$ as a function of $c$ must approach the value of $c$ where $\int_{S}\,(f-cg)=0$; call this value $c_{*}:=\int_{S}f/\int_{S}g$. More explicitly, $\|\tilde{u}\|_{\mathcal{H}}^{2}$ is unbounded as $\tilde{N},\tilde{N}_{\partial}\to\infty$ for $c=c_{*}\pm\varepsilon$ for any $\varepsilon>0$, but bounded at $c_{*}$, so the vertex of $\|\tilde{u}\|_{\mathcal{H}}^{2}$ as a function of $c$ must converge to $c_{*}$. This minimum (assuming $\boldsymbol{\Phi}$ is real) is at
(20) $\tilde{c}_{\tilde{N}}:=\frac{\boldsymbol{g}^{*}\boldsymbol{\Phi}^{-1}\boldsymbol{f}}{\boldsymbol{g}^{*}\boldsymbol{\Phi}^{-1}\boldsymbol{g}}\to c_{*},$
where we note that $\boldsymbol{\Phi}^{-1}$ is symmetric. Similar expressions to (20) appear in statistics literature when $\mathcal{H}$ is a Hilbert space of functions with zero average on $S$, such as Lemma 3 of [23] and Section 3.2 of [1]. We wish to work on more general surfaces where an exact basis for a space of functions that integrate to zero is not known, and therefore we use a basis of functions on a larger domain $\Omega$ instead (typically a box or all of $\mathbb{R}^{m}$). This means that a different approach was required to arrive at (20).
From (20), we conclude that as $\tilde{N},\tilde{N}_{\partial}\to\infty$,
@@
So, we have a linear operator to estimate $\int_{S}f$, with real weights as long as $\boldsymbol{\Phi}^{-1}$ is real (which is typical for a reasonable choice of $\mathcal{H}$). To be specific, let $\boldsymbol{w}^{*}=\int_{S}g\,\boldsymbol{\boldsymbol{g}}^{*}\boldsymbol{\Phi}^{-1}/(\boldsymbol{g}^{*}\boldsymbol{\Phi}^{-1}\boldsymbol{g})$, then $\boldsymbol{w}^{*}\boldsymbol{f}\to\int_{S}f$; $\boldsymbol{w}^{*}$ gives integration weights.
We may also want similar expressions to (20) and (21) in terms of inner products in $\mathcal{H}$. Let $v_{g}^{(\tilde{N})}$ and $v_{f}^{(\tilde{N})}$ be the solutions to
@@
where $F=g$ and $F=f$ to produce $v_{g}^{(\tilde{N})}$ and $v_{f}^{(\tilde{N})}$, respectively. These may correspond to problems on the domain $S$ that are not actually solvable, but the discretized version will still admit a unique solution. Then, as $\tilde{N},\tilde{N}_{\partial}\to\infty$,
@@ 3.1.1 Analysis
We now study the convergence rate of this method; in particular, we show that it inherits the convergence rate of the underlying meshfree PDE method.
Theorem 3.1.
Assume $u_{c}^{(\tilde{N})}$ is the solution to (19) for some values of $c$ and $\tilde{N}$. Assume that fcgf-cg is sufficiently smooth and $c$ is in any bounded interval, and, for some $p>0$, $\mathcal{H}$ produces a $p^{\text{th}}$-order meshfree method for problem (19) in the sense of (12):
@@
If an exact, extended, strong-form solution $u_{c_{*}}\in\mathcal{H}$ for $c_{*}=\int_{S}f/\int_{S}g$ exists for problem (17), then, where $\tilde{c}_{\tilde{N}}$ is the minimizer of $\big\|u_{c}^{(\tilde{N})}\big\|_{\mathcal{H}}$ as a function of $c$,
$\left|\tilde{c}_{\tilde{N}}\int_{S}g-\int_{S}f\right|\leq\tilde{A}_{p}h_{\text{max}}^{p}\left\|u_{c_{*}}\right\|_{\mathcal{H}}=\mathcal{O}\left(h_{\text{max}}^{p}\right),$
for sufficiently small $h_{\text{max}}$.
Before the proof of this theorem, it is important to note again here that $\big\|u_{c}^{(\tilde{N})}\big\|_{\mathcal{H}}$ may diverge as $\tilde{N}\to\infty$ and $h_{\text{max}}\to 0$ for a number of reasons; in fact it must diverge if (17) is not solvable for some choice of $c$ by Proposition 2.1. Also, the sufficient smoothness of fcgf-cg is only to allow the simplification of (12). If fcgf-cg, $S$, or $\partial S$ are such that (17) does not admit a solution that is smooth enough to be extended to $\mathcal{H}$, (22) and (23) still hold, but result in the divergence of $\big\|u_{c}^{(\tilde{N})}\big\|_{\mathcal{H}}$, rather than imply the convergence to zero of the left-hand sides of (22) and (23).
We now prove Theorem 3.1.
Proof 3.2.
Let $\left|\partial S\right|:=\int_{\partial S}1$, then (23) implies that
(24) $\left|\int_{S}\Delta_{S}u_{c}^{\left(\tilde{N}\right)}\right|=\left|\int_{\partial S}\hat{\boldsymbol{n}}_{\partial S}\cdot\nabla_{S}u_{c}^{\left(\tilde{N}\right)}\right|\leq\left|\partial S\right|B_{p}h_{\text{max}}^{p}\left\|u_{c}^{\left(\tilde{N}\right)}\right\|_{\mathcal{H}}$.
@@
In the case that $c=\int_{S}f/\int_{S}g$ and the PDE is solvable, this tells us nothing. However, this holds even when the PDE is not solvable, which gives us an estimate for the rate at which $\big\|u_{c}^{(\tilde{N})}\big\|_{\mathcal{H}}$ diverges:
@@
Now assume that an exact, extended, strong-form solution $u_{c_{*}}\in\mathcal{H}$ for $c_{*}=\int_{S}f/\int_{S}g$ exists for problem (17). In this case, $\big\|u_{c_{*}}^{(\tilde{N})}\big\|_{\mathcal{H}}\leq\left\|u_{c_{*}}\right\|_{\mathcal{H}}$ since $u_{c_{*}}$ is feasible for (19). Recall that $\tilde{c}_{\tilde{N}}=\left(\boldsymbol{g}^{*}\boldsymbol{\Phi}^{-1}\boldsymbol{f}\right)/\left(\boldsymbol{g}^{*}\boldsymbol{\Phi}^{-1}\boldsymbol{g}\right)$ is the minimizer of $\big\|u_{c_{*}}^{(\tilde{N})}\big\|_{\mathcal{H}}$ as a function of $c$ and that we already know that $\tilde{c}_{\tilde{N}}\to c_{*}$ from Eq. (20). Then,
@@
This estimate depends only on the norm of the PDE solution for $c=c_{*}$ and tells us that convergence of the integral estimate is at the rate at which $\Delta_{S}u_{c_{*}}^{(\tilde{N})}$ and $\hat{\boldsymbol{n}}_{\partial S}\cdot\nabla_{S}u$ converge when solving the PDE for $c=c_{*}$.